OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8 parallel ***perpendicular neither parallel nor perpendicular they are the same line

OpenStudy (the_fizicx99):

Substitute and solve, then graph it. 3x + 6(2x – 8) = 8 3x + 12x - 48 = 8 15x - 48 = 8 +48 +48 15x = 56 /15 /15 x = 3.73 .. So i can tell they're perpendicular